OpenStudy (anonymous):

Roots of the following function: g(x)=x^3+16x. Determine g(x)= a. x (x + 4) (x - 4) b. x (x + 4 i) ( x - 4 i) c. x ( x + 4) ( x + 4 i) d. x (x - 4) (x - 4 i)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Can you please show me how you got that answer?

OpenStudy (anonymous):

Ok. Hold on

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

x^3 + 16x = 0 Factor out the GCF, x. x(x^2 + 16) = 0 Set them equal to zero now. x = 0 x^2 + 16 = 0 x^2 = -16 x = ±√-16 x = ±i√16 x = ±4i Then, just bring it to the other side again. x(x + 4i)(x - 4i) = 0